# Long and Short Guidance in Score identity Distillation for One-Step Text-to-Image Generation

### v3

@@ Abstract
Diffusion-based text-to-image generation models trained on extensive text-image pairs have shown the capacity to generate photorealistic images consistent with textual descriptions. However, a significant limitation of these models is their slow sample generation, which requires iterative refinement through the same network. In this paper, we enhance Score identity Distillation (SiD) by developing long and short classifier-free guidance (LSG) to efficiently distill pretrained Stable Diffusion models without using real training data. SiD aims to optimize a model-based explicit score matching loss, utilizing a score-identity-based approximation alongside the proposed LSG for practical computation. By training exclusively with fake images synthesized with its one-step generator, SiD equipped with LSG rapidly improves FID and CLIP scores, achieving state-of-the-art FID performance while maintaining a competitive CLIP score. Specifically, its data-free distillation of Stable Diffusion 1.5 achieves a record low FID of 8.15 on the COCO-2014 validation set, with a CLIP score of 0.304 at an LSG scale of 1.5, and an FID of 9.56 with a CLIP score of 0.313 at an LSG scale of 2. Our code and distilled one-step text-to-image generators are available at https://github.com/mingyuanzhou/SiD-LSG.
@@ 1 Introduction
The pursuit of generating high-resolution, photorealistic images that matches the textual descriptions has consistently driven the machine learning community in developing powerful text-to-image generative models. Text-to-image diffusion models (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Podell et al., 2024) are currently leading the way in delivering unprecedentedly visual quality, diverse generation, and accurate text-image correspondences. They are renowned for their straightforward implementation and stability during optimization, and they receive substantial acclaim for the robust support from the open-source community (Rombach et al., 2022; Podell et al., 2024).
Despite these advantages, a significant limitation of diffusion models is their slow sampling process, which involves iterative refinement characterized by repeated forward passes through the generation network. Originally, this required a stochastic sampler with thousands of steps (Song and Ermon, 2019; Ho et al., 2020; Dhariwal and Nichol, 2021; Song et al., 2021a). Recent advancements in deterministic samplers based on ordinary differential equations have significantly reduced the required number of sampling steps to just tens or hundreds (Song et al., 2021a; Liu et al., 2022a; Lu et al., 2022b; Karras et al., 2022).
To further reduce the number of steps below ten or make a single step diffusion-based generator, the focus has shifted toward distilling the iterative-refinement based multi-step text-to-image generative progress into a single or a few steps, using a wide variety of acceleration techniques (Zheng et al., 2023b; Meng et al., 2023; Liu et al., 2022b; Luo et al., 2023b; Nguyen and Tran, 2024; Sauer et al., 2023b; Xu et al., 2023; Yin et al., 2023). However, they often result in clearly reduced ability to match the original data distribution, reflected as clearly worsening FIDs. All of them, with the exception of SwiftBrush (Nguyen and Tran, 2024), also require access to the original training data or the assistance of extra regression or adversarial losses.
@@
The success of SiD in distilling diffusion models in the pixel space has inspired us to apply it to open-source text-to-image latent diffusion models, specifically Stable Diffusion (SD) versions 1.5 and 2.1-base, aiming to significantly enhance their generation speed while maintaining performance. However, adapting SiD poses several notable challenges: first, SiD has primarily been applied to distill pre-trained EDM models, which utilize noise scheduling and preconditioning methods markedly different from the DDPM noise scheduling employed in SD; second, SiD-EDM does not incorporate classifier-free guidance (CFG) (Ho and Salimans, 2022), which is integral to SD; third, both the complexities and sizes of the data and model in EDM are much smaller than those in SD.
@@
To address these challenges, we explore the integration of CFG into SiD-based diffusion distillation. We have developed strategies that either enhance CFG on the pretrained score network, reduce it on the fake score network, or apply a combination of both, known as long and short guidance (LSG). This method allows SiD to efficiently distill SD models into one-step generators without needing training data. Remarkably, without relying on additional regression or adversarial losses—key components of recent distillation methods—or real data, our approach sets new benchmarks by achieving the lowest zero-shot FID score of 8.15 on the COCO-2014 (Lin et al., 2014) validation set. This significantly outperforms all previous distillation methods while maintaining competitive CLIP scores.
Figure 2: Rapid advancements in distilling Stable Diffusion 1.5 are showcased by the newly proposed SiD method, which employs long-short guidance (LSG). Key parameters include a batch size of 512, a learning rate of 1e-6, and an LSG scale of 2. This data-free approach achieves a zero-shot FID of 9.56 on the COCO-2014 validation set, along with a competitive CLIP score of 0.313. By reducing the LSG scale to 1.5, the FID can be further lowered to a record 8.15 among diffusion distillation models, with a corresponding CLIP score of 0.304. The series of images, generated from the same set of random noises post-training the SiD generator with varying counts of synthesized images, illustrates progressions at 0, 0.02, 0.1, 0.2, 0.5, 1, 2, 3, 4, and 5 million images. These are equivalent to 0, 40, 200, 400, 1k, 2k, 4K, 6K, 8k, and 10k training iterations respectively, organized from the top left to the bottom right. The progression of FIDs and CLIPs is detailed in the orange solid curves in the left plot of Fig. 4. The corresponding COCO-2014 validation text prompts are listed in Appendix E.
@@
The development of our SiD-LSG benefits tremendously from previous work on generative models, CFGs, and acceleration methods. We provide a detailed review of related work in Appendix B.
2 Empowering Score identity Distillation with long-short guidance
@@
The effectiveness of SiD has been demonstrated in distilling both unconditional and label-conditional EDM-based diffusion models provided by Karras et al. (2022), which are pretrained on pixel spaces up to a resolution of 64x64. In this paper, we expand the application of SiD to distill SD, the leading open-source platform for text-to-image diffusion models that operate on the latent space of an image encoder-decoder (Rombach et al., 2022). Our focus is specifically on SD-v1-5 (SD1.5) and SD-v2-1-base (SD2.1-base), which are two versions extensively benchmarked for diffusion distillation.
@@
The first challenge we tackle is adapting SiD, which was initially tested with EDM noise scheduling, to the classical DDPM scheduling employed by SD. A significant hurdle involves incorporating CFG (Ho and Salimans, 2022), which is essential for diffusion models to enhance their photorealism and text alignment, into the SiD loss functions. Addressing these challenges primarily requires modifying the derivation and loss functions of SiD. A third challenge arises because SD models are significantly larger and trained on data that is not only larger, but also more complex and of higher resolution. Overcoming this challenge requires addressing numerous technical details, such as establishing the minimum hardware requirements and computational precision. Additionally, it involves configuring the appropriate software settings to align with the constraints of the available computing platforms.
2.1 Semi-implicit distributions, score identities, and model-based score matching losses
@@
We denote ${\boldsymbol{c}}$ as the text representation of a pretrained text encoder, such as CLIP (Radford et al., 2021). Our objective is to distill a student model $p_{\theta}(\boldsymbol{x}_{g}\,|\,{\boldsymbol{c}})$ from a pretrained text-to-image diffusion model, such as SD1.5, which can generate text-guided random samples in a single step as: $\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z},{\boldsymbol{c}}),~{}\boldsymbol{z}\sim p(\boldsymbol{z}),$ where $G_{\theta}$ is a neural network parameterized by $\theta$ that deterministically transforms noise $\boldsymbol{z}\sim p(\boldsymbol{z})$ into generated data $\boldsymbol{x}_{g}$ under the guidance of text ${\boldsymbol{c}}$. The distribution of $\boldsymbol{x}_{g}$ is often implicit (Mohamed and Lakshminarayanan, 2016; Tran et al., 2017; Yin and Zhou, 2018), lacking an analytic probability density function (PDF) but is straightforward to sample from. The marginals of the real and generated data under the forward diffusion process can be expressed as:
@@
This structure, characterized by explicit conditional layers but implicit marginals, exemplifies a semi-implicit distribution (Yin and Zhou, 2018; Yu et al., 2023). This concept is employed by Zhou et al. (2024) to develop SiD, a method whose single-step data-free distillation capability has so far been demonstrated only on both unconditional and label-conditional diffusion models based on EDM.
We define the forward diffusion transition as $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\mathcal{N}(a_{t}\boldsymbol{x}_{0},\sigma_{t}^{2}\mathbf{I}),$ and unlike in SiD, which adheres to EDM noise scheduling where $a_{t}=1$, we allow $a_{t}$ to vary within $[0,1]$ to align with the DDPM scheduling used by SD during its training. This necessitates generalizing the equations used in SiD by permitting $a_{t}\neq 1$. We note that other diffusion types, such as categorical (Austin et al., 2021; Hoogeboom et al., 2021; Gu et al., 2022; Hu et al., 2022), Poisson (Chen and Zhou, 2023), and beta diffusions (Zhou et al., 2023), also align with the semi-implicit framework and can potentially be adapted similarly.
@@
MESM loss. A pretrained text-to-image diffusion model, such as SD, provides a score network $S_{\phi}$ parameterized by $\phi$ that estimates the true data score as
@@
It adopts $f_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ as the functional approximation of the conditional expectation of the real image $\boldsymbol{x}_{0}$ given noisy image $\boldsymbol{x}_{t}$ and text ${\boldsymbol{c}}$, expressed as $\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t},{\boldsymbol{c}}]$, adopts $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ to predict the noise inside $\boldsymbol{x}_{t}$, and adopts $-\sigma_{t}^{-1}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},{\boldsymbol{c}})$ as the functional approximation of the true score $\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}})$. Given time step $t\sim p(t)$ and text ${\boldsymbol{c}}$, we define the model-based explicit score-matching (MESM) distillation loss, which is a form of Fisher divergence (Lyu, 2009; Holmes and Walker, 2017; Yang et al., 2019; Yu and Zhang, 2023), as
@@
As $\psi^{*}(\theta)$ is unknown in practice, approximation is needed. We follow SiD to introduce an approximate optimization algorithm that alternates between optimize $\psi$ given $\theta$ using
@@
where $\boldsymbol{x}_{t}$ is generated as in (2) and $\omega(t)$ are weighted coefficients that will be specified. Note that (5) simplifies the corresponding one in SiD by fixing $\alpha=1$, a parameter that can be tuned in SiD to ehance performance but fixed at one in this paper. We have strategically set $\alpha=1$ to focus more on analyzing the injection of CFGs into SiD, leaving the tuning of this parameter for future work.
@@ 2.2 SiD with classifier-free guidance
An essential practice for enhancing photorealism and alignment with text instructions in text-to-image diffusion models involves incorporating CFG into reverse diffusion. This principle also applies to distillation methods for these models, where CFG must be integrated into the appropriate terms of their distillation loss functions. Therefore, a key distinction in the distillation of SD, compared to previous unconditional and label-conditional diffusion models, lies in the need to introduce CFG.
SiD presents a unique opportunity to apply CFG to enhance its text-to-image generation performance. First, we note CFG enhances text guidance by modifying the distribution of $\boldsymbol{x}_{t}$ given text ${\boldsymbol{c}}$ as
$p(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}},\kappa)\propto p({\boldsymbol{c}}\,|\,\boldsymbol{x}_{t})^{\kappa}p(\boldsymbol{x}_{t})\propto\left(\frac{p(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}})}{p(\boldsymbol{x}_{t})}\right)^{\kappa}p(\boldsymbol{x}_{t}),$
@@
  Input: Pretrained score network $f_{\phi}$, generator $G_{\theta}$, fake score network $f_{\psi}$, $t_{\text{init}}=625$, $t_{\min}=20$, $t_{\max}=979$, guidance scales $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=2$
  Initialization $\theta\leftarrow\phi,\psi\leftarrow\phi$
  repeat
    Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and ${\boldsymbol{c}}$, and let $\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z},{\boldsymbol{c}})=f_{\theta}(\sigma_{t_{\text{init}}}\boldsymbol{z},t_{\text{init}},{\boldsymbol{c}})$; Sample $t\in\{t_{min},\ldots,t_{\max}\}$ and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$, and let $\boldsymbol{x}_{t}=a_{t}\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$
  until processing 10M fake images or the training budget is exhausted
  Output: $G_{\theta}$

### v4

# Guided Score identity Distillation for Data-Free One-Step Text-to-Image Generation

@@ Abstract
Diffusion-based text-to-image generation models trained on extensive text-image pairs have demonstrated the ability to produce photorealistic images aligned with textual descriptions. However, a significant limitation of these models is their slow sample generation process, which requires iterative refinement through the same network. To overcome this, we introduce a data-free guided distillation method that enables the efficient distillation of pretrained Stable Diffusion models without access to the real training data, often restricted due to legal, privacy, or cost concerns. This method enhances Score identity Distillation (SiD) with Long and Short Classifier-Free Guidance (LSG), an innovative strategy that applies Classifier-Free Guidance (CFG) not only to the evaluation of the pretrained diffusion model but also to the training and evaluation of the fake score network. We optimize a model-based explicit score matching loss using a score-identity-based approximation alongside our proposed guidance strategies for practical computation. By exclusively training with synthetic images generated by its one-step generator, our data-free distillation method rapidly improves FID and CLIP scores, achieving state-of-the-art FID performance while maintaining a competitive CLIP score. Notably, the one-step distillation of Stable Diffusion 1.5 achieves an FID of 8.15 on the COCO-2014 validation set, a record low value under the data-free setting. Our code and checkpoints are available at https://github.com/mingyuanzhou/SiD-LSG.
@@ 1 Introduction
The pursuit of generating high-resolution, photorealistic images that matches the textual descriptions has driven the machine learning community in developing powerful text-to-image (T2I) generative models. T2I diffusion models (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Podell et al., 2024) are currently leading the way in delivering unprecedentedly visual quality, diverse generation, and accurate text-image correspondences. They are renowned for their straightforward implementation and stability during optimization, and they receive substantial acclaim for the robust support from the open-source community (Rombach et al., 2022; Podell et al., 2024).
Despite these advantages, a significant limitation of diffusion models, including those used for T2I tasks, is their slow sampling process, which involves iterative refinement through repeated passes of the generation network. Originally, this required thousands of stochastic sampling steps (Song & Ermon, 2019; Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021a). Recent advancements in ODE-based deterministic samplers have significantly reduced the required number of sampling steps to just tens or hundreds (Song et al., 2021a; Liu et al., 2022a; Lu et al., 2022b; Karras et al., 2022). To further reduce the number of steps below ten, or even down to one, the focus has shifted toward distilling the iterative-refinement based multi-step T2I generative progress, using a wide variety of acceleration techniques (Zheng et al., 2023b; Meng et al., 2023; Liu et al., 2022b; Luo et al., 2023b; Nguyen & Tran, 2024; Sauer et al., 2023b; Xu et al., 2023; Yin et al., 2023). However, they often result in clearly reduced ability to match the original data distribution, reflected as clearly worsening FIDs. All of them, with the exception of SwiftBrush (Nguyen & Tran, 2024), also require access to real images or the assistance of extra regression or adversarial losses.
@@
The success of SiD in distilling EDM diffusion models for non-T2I generation in the pixel space has inspired us to adapt it to open-source T2I latent diffusion models, specifically Stable Diffusion (SD) versions 1.5 and 2.1-base, aiming to significantly enhance their generation speed while maintaining performance. However, adapting SiD poses several notable challenges: first, SiD-EDM does not incorporate classifier-free guidance (CFG) (Ho & Salimans, 2022), which is integral to SD; second, SiD has primarily been applied to distill pre-trained EDM models, which utilize noise scheduling and preconditioning methods markedly different from the DDPM noise scheduling employed in SD; third, both the complexities and sizes of the data and model in EDM are much smaller than those in SD.
@@
To address these challenges, we explore the integration of CFG and SiD for T2I diffusion distillation. In addition to testing the conventional approach of enhancing CFG on the pretrained score network, we introduce a novel strategy of reducing CFG on the fake score network, as well as a combined approach that employs both strategies, which we refer to as long and short guidance (LSG). This new method efficiently distills SD models into one-step T2I generators without requiring training data. Surprisingly, it achieves a new benchmark on the COCO-2014 validation set with a zero-shot FID score of 8.15, the lowest to date for one-step, data-free distillation, despite not relying on additional regression or adversarial losses, nor real data—key components of recent distillation techniques.
@@
The development of our SiD-LSG approach builds significantly on prior work in generative models, CFG, and acceleration methods. A comprehensive review of related work is provided in Appendix B.
Figure 2: Rapid advancements in distilling Stable Diffusion 1.5 are showcased by the proposed SiD method that incorporates long-short guidance (LSG). Key parameters include a batch size of 512, a learning rate of 1e-6, and an LSG scale of 2. This data-free approach achieves a zero-shot FID of 9.56 on the COCO-2014 validation set, along with a competitive CLIP score of 0.313. By reducing the LSG scale to 1.5, the FID can be further lowered to a record 8.15 among data-free diffusion distillation models, with a corresponding CLIP score of 0.304. The series of images, generated from the same set of random noises post-training the SiD generator with varying counts of synthesized images, illustrates progressions at 0, 0.02, 0.1, 0.2, 0.5, 1, 2, 3, 4, and 5 million images. These are equivalent to 0, 40, 200, 400, 1k, 2k, 4K, 6K, 8k, and 10k training iterations respectively, organized from the top left to the bottom right. The progression of FIDs and CLIPs is detailed in the orange solid curves in the left plot of Fig. 4. The corresponding COCO-2014 validation text prompts are listed in Appendix F.
2 Data-Free Guided Score identity Distillation
@@
We explore the use of SiD and CFG to distill SD, the leading open-source platform for T2I diffusion models that operate on the latent space of an image encoder-decoder (Rombach et al., 2022), with a specific focus on the data-free setting. Our focus is specifically on SD-v1-5 (SD1.5) and SD-v2-1-base (SD2.1-base), which are two versions extensively benchmarked for diffusion distillation.
@@
A significant hurdle involves incorporating CFG (Ho & Salimans, 2022), essential for T2I diffusion models to enhance their photorealism and text alignment, into the SiD loss functions. The second challenge is adapting SiD, initially tested with EDM noise scheduling, to the DDPM scheduling employed by SD. Addressing these challenges primarily requires modifying the derivation and loss functions of SiD. A third challenge arises because SD models are significantly larger and trained on bigger, more complex, and higher resolution data. Overcoming this challenge requires addressing numerous technical details, such as establishing the minimum hardware requirements and configuring the appropriate software settings to align with the constraints of the available computing platforms.
2.1 Preliminaries on Score identity Distillation
@@
We denote ${\boldsymbol{c}}$ as the text representation of a pretrained text encoder, such as CLIP (Radford et al., 2021). Our objective is to distill a student model $p_{\theta}(\boldsymbol{x}_{g}\,|\,{\boldsymbol{c}})$ from a pretrained T2I diffusion model, such as SD1.5, which can generate text-guided random samples in a single step as: $\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z},{\boldsymbol{c}}),~{}\boldsymbol{z}\sim p(\boldsymbol{z}),$ where $G_{\theta}$ is a neural network parameterized by $\theta$ that deterministically transforms noise $\boldsymbol{z}\sim p(\boldsymbol{z})$ into generated data $\boldsymbol{x}_{g}$ under the guidance of text ${\boldsymbol{c}}$. The distribution of $\boldsymbol{x}_{g}$ is often implicit (Mohamed & Lakshminarayanan, 2016; Tran et al., 2017; Yin & Zhou, 2018), lacking an analytic probability density function (PDF) but is straightforward to sample from. The marginals of the real and generated data under the forward diffusion process can be expressed as:
@@
This structure, characterized by explicit conditional layers but implicit marginals, exemplifies a semi-implicit distribution (Yin & Zhou, 2018; Yu et al., 2023). This concept is employed by Zhou et al. (2024) to develop SiD, a method whose single-step data-free distillation capability has so far been demonstrated only on non-T2I diffusion models based on EDM.
We define the forward diffusion transition as $q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{0})=\mathcal{N}(a_{t}\boldsymbol{x}_{0},\sigma_{t}^{2}\mathbf{I}),$ and unlike in SiD, which adheres to EDM noise scheduling where $a_{t}=1$, we allow $a_{t}$ to vary within $[0,1]$ to align with the DDPM scheduling used by SD. This necessitates generalizing the equations used in SiD by permitting $a_{t}\neq 1$. We note that other diffusion types, such as categorical (Austin et al., 2021; Hoogeboom et al., 2021; Gu et al., 2022; Hu et al., 2022), Poisson (Chen & Zhou, 2023), and beta diffusions (Zhou et al., 2023), also align with the semi-implicit framework and can potentially be adapted similarly.
@@
MESM loss. A pretrained T2I diffusion model, such as SD, provides a score network $S_{\phi}$ parameterized by $\phi$ that estimates the true data score as
@@
It adopts $f_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ as the functional approximation of the conditional expectation of the real image $\boldsymbol{x}_{0}$ given noisy image $\boldsymbol{x}_{t}$ and text ${\boldsymbol{c}}$, expressed as $\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t},{\boldsymbol{c}}]$, adopts $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ to predict the noise inside $\boldsymbol{x}_{t}$, and adopts $-\sigma_{t}^{-1}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},{\boldsymbol{c}})$ as the functional approximation of the true score $\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}})$. Given time step $t\sim p(t)$ and text ${\boldsymbol{c}}$, we define the model-based explicit score-matching (MESM) distillation loss, which is a form of Fisher divergence (Lyu, 2009; Holmes & Walker, 2017; Yang et al., 2019; Yu & Zhang, 2023), as
@@
As $\psi^{*}(\theta)$ is unknown in practice, we follow SiD to alternates between optimizing $\psi$ and $\theta$ using
@@
where $\boldsymbol{x}_{t}$ is generated as in (2) and $\omega(t)$ are weighted coefficients that will be specified.
@@ 2.2 SiD with classifier-free guidance
An essential practice for enhancing photorealism and alignment with text instructions in T2I diffusion models involves incorporating CFG into reverse diffusion. This principle also applies to distillation methods for these models, where CFG must be integrated into the appropriate terms of their distillation loss functions. Therefore, a key distinction in the distillation of SD, compared to previous unconditional and label-conditional diffusion models, lies in the need to introduce CFG.
SiD presents a unique opportunity to apply CFG to enhance its T2I generation performance. First, we note CFG enhances text guidance by modifying the distribution of $\boldsymbol{x}_{t}$ given text ${\boldsymbol{c}}$ as
$p(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}},\kappa)\propto p({\boldsymbol{c}}\,|\,\boldsymbol{x}_{t})^{\kappa}p(\boldsymbol{x}_{t})\propto\big{(}\frac{p(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}})}{p(\boldsymbol{x}_{t})}\big{)}^{\kappa}p(\boldsymbol{x}_{t}),$